# Benchmarking direct and indirect dipolar spin-exchange interactions between two Rydberg atoms

### v1

@@ Appendix A Effect of positional disorder onto the damping of a spin exchange oscillation
$\langle\mathbf{r}\ket{\psi_{\mathrm{rel}}(t)}=\varphi_{\uparrow,\downarrow}(\mathbf{r},t)\ket{\uparrow,\downarrow}+\varphi_{\downarrow,\uparrow}(\mathbf{r},t)\ket{\downarrow,\uparrow}$ (4)
@@
$i\hbar\frac{\partial\ket{\psi_{\mathrm{rel}}}}{\partial t}=H(\hat{\mathbf{r}})\ket{\psi_{\mathrm{rel}}}.$ (5)
where $\hat{\mathbf{r}}$ is the distance operator, and the Hamiltonian is given by
$H(\hat{\mathbf{r}})=\hbar\begin{bmatrix}0&J(\hat{\mathbf{r}})\\ J(\hat{\mathbf{r}})&0\end{bmatrix}$ (6)
in the basis {$\{\ket{\uparrow,\downarrow},\ket{\downarrow,\uparrow}\}$. Neglecting the angular dependence, the spin exchange term can be written $J(\hat{\mathbf{r}})\approx J(\hat{r})=J_{0}\left(r_{0}/\hat{r}\right)^{\alpha}$ with $r_{0}=|\mathbf{r}_{0}|$ the average distance between the atoms and $\alpha=3$ (resp. $\alpha=6$) for a direct (resp. indirect) spin-exchange. Note that compared with Hamiltonian (2) of the main text, we have neglected the diagonal van der Waals term, which will only lead to a global phase accumulation. Defining $\varphi_{\pm}\equiv\frac{1}{\sqrt{2}}\left(\varphi_{\uparrow,\downarrow}\pm\varphi_{\downarrow,\uparrow}\right)$, Eq. (5) can be rewritten as
$\displaystyle\frac{\partial\varphi_{\pm}}{\partial t}(\mathbf{r},t)=\mp iJ(r)\varphi_{\pm}(\mathbf{r},t).$ (7)
In a spin-exchange experiment such as described in Fig. 2(a), the initial state is given by the product state
$\langle\mathbf{r}\ket{\psi_{\mathrm{rel}}(t=0)}$ $=\varphi_{0}(\mathbf{r})\ket{\downarrow,\uparrow}$
$=\frac{\varphi_{0}(\mathbf{r})}{\sqrt{2}}\left(\ket{+}+\ket{-}\right)$ (8)
with $|\varphi_{0}(\mathbf{r})|^{2}$ the initial distribution of interatomic distances, centered around $\mathbf{r}_{0}$. Neglecting the modification of the spatial wave function due to the force induced by the interaction the force, the resulting time evolution following Eq. (7) leads to a space-dependent phase accumulation for the wavefunction:
$\langle\mathbf{r}\ket{\psi_{\mathrm{rel}}(t)}=\frac{\varphi_{0}(\mathbf{r})}{\sqrt{2}}\left(e^{-iJ(r)t}\ket{+}+e^{iJ(r)t}\ket{-}\right).$ (9)
Our experiment measures the probability of the state $\ket{\uparrow,\downarrow}$ with no access to the position information, that is to say
$P_{\uparrow,\downarrow}(t)$ $=\int d\mathbf{r}^{3}\left|\langle\mathbf{r},\uparrow,\downarrow\ket{\psi_{\mathrm{rel}}(t)}\right|^{2}$
@@
This formula shows that even for trapped Rydberg atoms without any dipole-dipole force, the spin-exchange oscillation is fundamentally limited by the spread of fluctuations of the distribution of distances. For perfectly-localized atoms ($|\varphi_{0}(\mathbf{r})|^{2}=\delta^{(3)}(\mathbf{r}-\mathbf{r}_{0})$), one recovers a perfect sinusoidal oscillation with frequency $2J_{0}$: $P_{\uparrow,\downarrow}(t)=\cos^{2}(J_{0}t)$.
To push the calculation further, we simplify the problem to a single radial dimension $x$ and consider a Gaussian distribution of positions: $|\varphi_{0}(\mathbf{r})|^{2}=\delta(y)\,\delta(z)\,\exp\left[-(x-r_{0})^{2}/2\Delta r^{2}\right]/\sqrt{2\pi}\Delta r$, where the standard deviation $\Delta r$ relates to the single-atom radial dispersion by $\Delta r=\sqrt{2}\sigma_{\perp}+O(\sigma_{\perp}^{2}/r)$. Given that $\Delta r\ll r_{0}$, we linearize $J(r)$ around $r_{0}$: $J(r)\approx J_{0}-\alpha J_{0}(r-r_{0})/r_{0}$. In this case, the Gaussian distribution of distances translates into a Gaussian distribution of interaction energies with standard deviation $\Delta J=\alpha J_{0}\Delta r/r_{0}$, leading to
@@
$\approx\frac{1}{2}\left[1+e^{-2\Delta J^{2}t^{2}}\cos(2J_{0}t)\right]$ (11)
The spin-exchange oscillation is damped by a Gaussian envelope $e^{-t^{2}/2\tau^{2}}$, with a damping time $\tau=1/2\Delta J$. This corresponds to a quality factor $Q=\tau J_{0}=r_{0}/2\alpha\Delta r$.
Appendix B Condition to neglect the dipole-dipole forces
Assuming that the atoms are in free flight amounts to neglecting the forces induced by the dipole-dipole interaction. This assumption is only valid at short times, when the dipole-dipole force has not yet affected the atomic positions. The time at which this assumption breaks down can be estimated from classical arguments. In the case of Fig. 2(b), the dominant potential is $\hbar J_{SP}(r)$, which gives rise to a force
$F(r)$ $=-\hbar\frac{\partial J_{SP}}{\partial r}(r)$
$=\frac{3\hbar J_{SP}(r_{0})r_{0}^{3}}{r^{4}}$ (12)
The equation of motion for the interatomic distance is
$\mu\frac{d^{2}r}{dt^{2}}=F(r)$ (13)
with $\mu=m/2$ the reduced mass of the two-atom system. If we approximate the force by its value at the initial position $r_{0}$ ($F(r)\sim F(r_{0})$), we obtain the displacement
$r(t)-r_{0}\sim\frac{3\hbar J_{SP}(r_{0})}{2\mu r_{0}}t^{2}.$ (14)
The time at which this displacement gets on the same order of magnitude as the initial positional disorder ($\Delta r\sim\sqrt{2}\sigma_{\perp}$) is given by
$r(t)-r_{0}\sim\sqrt{2}\sigma_{\perp}\iff t\sim\sqrt{\frac{2\sqrt{2}\mu r_{0}\sigma_{\perp}}{3\hbar J_{SP}(r_{0})}}$ (15)
With the parameters of Fig. 2, we find $t\sim 14$ µs, which is indeed longer than our experimental durations.

### v2

@@ Appendix B A simple analytic model for the damping of a spin exchange oscillation
$r(t,x_{0},$ $v_{0})=$ (8)
$r_{0}\;+$$\left\{\begin{array}[]{ll}x_{0}+v_{0}t\quad\text{for atoms in free flight}\\[6.0pt] x_{0}\cos(\omega_{\perp}t)+\frac{v_{0}}{\omega_{\perp}}\sin(\omega_{\perp}t)\quad\text{for trapped atoms}.\end{array}\right.$ (11)
with $r_{0}$ the average distance.
Each realization of the positional disorder $(x_{0},v_{0})$ corresponds to a given dynamics for the internal state |$\ket{\psi(t,x_{0},v_{0})}$, which is governed by the Schrödinger equation:
$i\hbar\frac{d\ket{\psi}}{dt}=H\left[r(t,x_{0},v_{0})\right]\ket{\psi}\ ,$ (12)
where the Hamiltonian for the internal state is given by
$H(r)=\hbar\begin{bmatrix}0&J(r)\\ J(r)&0\end{bmatrix}$ (13)
in the basis {$\{\ket{\uparrow,\downarrow},\ket{\downarrow,\uparrow}\}$. The spin exchange coupling can be written $J(r)=J_{0}\left(r_{0}/r\right)^{\alpha}$ with $\alpha=3$ (resp. $\alpha=6$) for a direct (resp. indirect) spin-exchange. Note that compared with Hamiltonian (2) of the main text, we have neglected the diagonal van der Waals term, which will only lead to a global phase accumulation. Writing |$\ket{\psi}=c_{+}\ket{+}+c_{-}\ket{-}$ with |±=(|,±|$\ket{\pm}=(\ket{\uparrow,\downarrow}\pm\ket{\downarrow,\uparrow})/\sqrt{2}$, Eq. (12) becomes
$\frac{dc_{\pm}}{dt}=\mp iJ\left[r(t,x_{0},v_{0})\right]c_{\pm}$ (14)
whose solution is
@@
for the initial state of a spin exchange |$\ket{\psi(t=0)}=\ket{\uparrow,\downarrow}$. Each realization of the positional disorder leads to the following probability for the state $\ket{\uparrow,\downarrow}$:
$p_{\uparrow,\downarrow}(t,x_{0},v_{0})$ =|,$=\left|\langle\uparrow,\downarrow\ket{\psi(t,x_{0},v_{0})}\right|^{2}$ (16)
$=\frac{1}{2}\left|c_{+}(t,x_{0},v_{0})+c_{-}(t,x_{0},v_{0})\right|^{2}$
@@
Finally, our experiment measures the probability of the state $\ket{\uparrow,\downarrow}$ averaged over all realizations of the positional disorder:
@@
We now evaluate Eq. (17) in the case of (i) atoms in free flight and (ii) trapped atoms, following the classical trajectories given by Eq. (8). To do so, we linearize $J(r)$ around $r_{0}$: $J(r)\approx J_{0}-\alpha J_{0}(r-r_{0})/r_{0}$ which is valid if $|r-r_{0}|\ll r_{0}$. The Gaussian distribution of distances translates into a Gaussian distribution of interaction energies with standard deviation $\Delta J=\alpha J_{0}\Delta r/r_{0}$, allowing us to use Gaussian integration formulas.
(i) For atoms in free flight, we obtain
$P_{\uparrow,\downarrow}^{\text{free}}(t)=\frac{1}{2}\left[1+e^{-2\Delta J^{2}t^{2}\,\left(1+\frac{1}{4}\omega_{\perp}^{2}t^{2}\right)}\cos\left(2J_{0}t\right)\right].$ (18)
The amplitude of the oscillation is reduced by an exponential envelope which contains a quadratic term $\propto t^{2}$ due to the initial position dispersion, and a quartic term $\propto t^{4}$ which originates from the initial velocity dispersion. In Fig. 6(a, top panel), we compare the theory profile of Eq. (18) to a numerical simulation for the experimental parameters of Fig. 2(c), assuming that atoms are perfectly cooled to the motional ground state. We obtain a perfect agreement, thus validating the approximations which were made in the derivation, namely: neglecting positional disorder along $z$ and linearizing the dipole-dipole potential. We also include finite Rydberg lifetimes in the simulation (bottom panel) and conclude that positional disorder is the main limitation for atoms in free flight in this range of parameters.
The $1/\sqrt{e}$ damping time in Eq. (18) is given by
$\tau$ $=\frac{\sqrt{2}}{\omega_{\perp}}\sqrt{\sqrt{1+\frac{\omega_{\perp}^{2}}{4\Delta J^{2}}}-1}$ (19)
The dependence of the quality factor $Q=J_{0}\tau$ on the initial trapping frequency $\omega_{\perp}$ is plotted as a dashed blue line in Fig. 6(c) for the parameters of Fig. 2(c) and for $\bar{n}_{\perp}=0$. Using the fact that $\Delta J$ depends on $\omega_{\perp}$ as $\Delta J=\alpha J_{0}\Delta r(\omega_{\perp})/r_{0}$, we find that the quality factor is maximized for an optimal trapping frequency
$\omega_{\perp}^{\mathrm{opt}}=2\left[\frac{4\alpha^{2}\hbar(2\bar{n}_{\perp}+1)J_{0}^{2}}{mr_{0}^{2}}\right]^{1/3},$ (20)
corresponding to $\tau(\omega_{\perp}^{\mathrm{opt}})=2/\omega_{\perp}^{\mathrm{opt}}$. We find $\omega_{\perp}^{\mathrm{opt}}=2\pi\times 36$ kHz and $Q(\omega_{\perp}^{\mathrm{opt}})=34$. The dominant contribution to the damping is the initial position dispersion for $\omega_{\perp}<\omega_{\perp}^{\mathrm{opt}}$ and the initial velocity dispersion for $\omega_{\perp}>\omega_{\perp}^{\mathrm{opt}}$. Numerical simulations in the presence of finite Rydberg lifetimes [solid blue lines in Fig. 6(c)] show that the quality factor is weakly affected by the lifetimes in this range of parameters.
(ii) For trapped atoms, we obtain
$P_{\uparrow,\downarrow}^{\text{trapped}}(t)=\frac{1}{2}\left[1+e^{-4\left(\frac{\Delta J}{\omega_{\perp}}\right)^{2}\left(1-\cos(\omega_{\perp}t)\right)}\cos\left(2J_{0}t\right)\right].$ (21)
The envelope shows periodic revivals due to the oscillations of the atomic positions in the traps, which cause the probability of each classical trajectory [Eq. (16)] to rephase. Those revivals were also derived in Ref. [48] using a full quantum treatment of the atomic motion (see Appendix C). The modulation of the envelope in Eq. (21) can be suppressed by increasing the trapping frequency $\omega_{\perp}$ so to reach the regime $\omega_{\perp}\gg\Delta J$. Using the dependence of $\Delta J$ on $\omega_{\perp}$, the condition to suppress the modulation of the envelope becomes $\omega_{\perp}\gg\omega_{\perp}^{\mathrm{crit}}$ with a critical trapping frequency
$\omega_{\perp}^{\mathrm{crit}}=\left[\frac{\alpha J_{0}}{r_{0}}\sqrt{\frac{\hbar(2\bar{n}_{\perp}+1)}{m}}\right]^{2/3}.$ (22)
For the parameters of Fig. 2(c), we find $\omega_{\perp}^{\mathrm{crit}}=2\pi\times 11$ kHz in the ground state.
In Fig. 6(b, top panel), we simulate the spin exchange of two trapped atoms in their motional ground state for $\omega_{\perp}=2\pi\times 37$ kHz $>\omega_{\perp}^{\mathrm{crit}}$ and find again a good agreement with the predicted profile of Eq. (21). Now finite Rydberg lifetimes become the dominating limiting factor to the damping [Fig. 6(b, bottom panel)], leading to a significant improvement in the coherence of the oscillation compared with the case of free atoms.
To estimate by how much Rydberg trapping can improve the damping, we repeat the simulations for various initial trapping frequencies $\omega_{\perp}$ and extract the quality factor of each resulting spin-exchange oscillation. Due to the revivals, fitting $P_{\uparrow\downarrow}(t)$ by a Gaussian envelope is not reliable, so we define the damping time $\tau$ as the $1/\sqrt{e}$ decay time for the contrast of $P_{\uparrow\downarrow}(t)$. The results are shown in Fig. 6(c) as orange lines: the quality factor reaches $\sim 100$ for trapped Rydberg atoms and saturates starting from $\omega_{\perp}\sim 2\pi\times 37$ kHz due to the finite lifetimes.
Appendix C Quantum treatment of positional disorder
Numerical simulations performed throughout the paper are Monte-Carlo simulations where atomic positions and velocities are sampled from Gaussian distributions and evolved according to classical trajectories (similarly to Refs. [7, 10, 9, 8]). One may wonder whether the results remain valid as the atomic wavefunctions are approaching the ground state of the trap. This case is treated in details in Ref. [48, 12], here we simply implement numerical simulations of the quantum equations of motion with our experimental parameters.
Restricting the internal degrees of freedom to $\ket{\uparrow,\downarrow}$ and $\ket{\downarrow,\uparrow}$, the two-atom system is described by the following relative wavefunction:
$\langle r\ket{\psi_{\mathrm{rel}}(t)}=\varphi_{\uparrow,\downarrow}(r,t)\ket{\uparrow,\downarrow}+\varphi_{\downarrow,\uparrow}(r,t)\ket{\downarrow,\uparrow}$ (23)
[…]
$i\hbar\frac{d\ket{\psi_{\mathrm{rel}}}}{dt}=\left(\frac{\hat{p}^{2}}{2\mu}+\frac{1}{2}\mu\omega_{\perp}^{2}(\hat{r}-r_{0})^{2}+H(\hat{r})\right)\ket{\psi_{\mathrm{rel}}}.$ (24)
where $\hat{r}$ is the distance operator, $r_{0}$ is the average distance, $\mu=m/2$ is the reduced mass of the two-atom system, and the effective Hamiltonian for the dipole-dipole interactions is given by
$H(\hat{r})=\hbar\begin{bmatrix}V(\hat{r})&J(\hat{r})\\ J(\hat{r})&V(\hat{r})\end{bmatrix}$ (25)
in the basis {$\{\ket{\uparrow,\downarrow},\ket{\downarrow,\uparrow}\}$. Compared to Hamiltonian (13), we include the diagonal term $V$ as it can affect the motion of the wavepackets. Equation (24) can be rewritten for $\varphi_{\pm}=\left(\varphi_{\uparrow,\downarrow}\pm\varphi_{\downarrow,\uparrow}\right)/\sqrt{2}$ as two independent differential equations:
$\displaystyle\frac{\partial\varphi_{\pm}}{\partial t}=\displaystyle\frac{i\hbar}{2\mu}\frac{\partial^{2}\varphi_{\pm}}{\partial r^{2}}-iV_{\pm}(r)\varphi_{\pm}$ (26)
where we have defined the state-dependent potential $V_{\pm}(r)=V(r)\pm J(r)+\frac{1}{2}\mu\omega_{\perp}^{2}(r-r_{0})^{2}$. An important consequence of this potential is the possibility of spin-motion coupling, meaning that atomic motion is affected by the internal state — an effect which is not taken into account in the classical Monte-Carlo simulation.
We perform numerical simulations of the quantum equations of motion (26) using a Crank–Nicolson scheme, starting from the state $\langle r\ket{\psi_{\mathrm{rel}}(t=0)}=\varphi_{0}(r)\ket{\downarrow,\uparrow}$ with $|\varphi_{0}(r)|^{2}=\exp\left[-(r-r_{0})^{2}/2\Delta r^{2}\right]/\sqrt{2\pi}\Delta r$ the distribution of interatomic distances in the ground state of the trap at frequency $\omega_{\perp}$. We then evaluate the probability $P_{\uparrow\downarrow}(t)=\int dr|\langle r,\uparrow,\downarrow\ket{\psi_{\mathrm{rel}}(t)}|^{2}$. We also add the finite Rydberg lifetimes as an additional decoherence channel using a quantum Monte-Carlo method [51].
The results are shown as dotted lines in Fig. 6(c). We find that they give the same damping as the classical Monte-Carlo simulations —we also checked that the probabilities $P_{\uparrow\downarrow}(t)$ from both simulation methods overlap perfectly in the time domain—, confirming that for our range of parameters, quantum fluctuations of positions and velocities can be safely replaced by classical fluctuations, and that spin-motion coupling does not affect the damping of the oscillation.